a car moving at 10m/s accelerated uniformly through a distance of 25m/s at the end which the velocity of the car became 15m/s.
determine the acceleration of the car and the time taken to cover 25m
1
Expert's answer
2019-09-26T14:48:42-0400

The distance traveled by car


"d=\frac{v_f^2-v_i^2}{2a}"

So, acceleration of the car


"a=\frac{v_f^2-v_i^2}{2d}=\frac{15^2-10^2}{2\times 25}=2.5\:\rm m/s^2"

Time of motion

"t=\frac{v_f-v_i}{a}=\frac{15-10}{2.5}=2\:\rm s"